OpenStudy (anonymous):

Which is NOT rational? (As in rational number) 4 + sqrt(2^2 -1) 1/3 sqrt(6^2 + 13) 9

OpenStudy (anonymous):

1/3 I think.

OpenStudy (anonymous):

4/5 is a rational number so i do believe that 1/3 is also

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

a rational number is a number that, when written in decimal form, either repeats endlessly, OR has a decimal that ends. 4/5 = 0.8, so it is a rational number becuase the decimal ends 1/3 = 0.33333333333333333, is a rational number because it repeats endlessly make sense now?

OpenStudy (anonymous):

but in my lesson for class it said that 2.828427125 was irrational and -3.141592654.. was irrational also is it because there are a bunch of different numbers after the decimal point instead of just one?

OpenStudy (anonymous):

my lessons definition was A rational number is any number that can be written as a fraction, or a ratio of two integers.

OpenStudy (anonymous):

OOOOOHHHH im sorry i misunderstood everything that you just said.

OpenStudy (anonymous):

I know both fractions are rational its just the other two options xD

OpenStudy (anonymous):

"but in my lesson for class it said that 2.828427125 was irrational and -3.141592654.. was irrational also is it because there are a bunch of different numbers after the decimal point instead of just one?" very good, you get it. an irrational number is when it just has random numbers after it, yep

OpenStudy (anonymous):

4+ sqrt(2^2 - 1) 4 - 1 is 3 the square root of 3 is 1.73 1.73 + 4 = 5.73 so that is rational?

OpenStudy (anonymous):

is the sqrt(3) = 1.73? or does it actully continue randomlyh?

OpenStudy (anonymous):

oh, its a bunch of random numbers xD

OpenStudy (anonymous):

very good. a bunch of random numbers means Irrational

OpenStudy (anonymous):

then the other one would be 49 which would equal 7 and yeah that ones irrational

OpenStudy (anonymous):

Alright I got it now, thanks xD

OpenStudy (anonymous):

^_^